# Toward Robust Imperceptible Perturbation against Unauthorized Text-to-image Diffusion-based Synthesis

### v2

@@ 1 Introduction
Diffusion models achieve significant success in a wide range of applications, including image generation [15; 47; 7], image editing [21; 46; 5], and text-to-image synthesis [38]. Subject-driven text-to-image synthesis, an emerging application of diffusion models, in particular, has attracted considerable attention due to its potential to generate personalized images from a few reference photos. Among the approaches proposed to achieve this goal [34; 42], DreamBooth [40] and Text Inversion [12] are two prominent training-based methods that offer impressive personalize generation ability. With an additional lightweight model fine-tuning process for capturing the subject or a concept-related embedding training phase, personalized diffusion models can retain the generation capacity from the pre-training stage and conduct vivid personalized generation. While these methods empower high-quality personalized generation, they also raise privacy concerns as they can fabricate misleading or harmful content in the wrong hands, endangering individuals. For example, recent news [19] indicates that AI tools like diffusion models have been employed to generate fake profiles of individuals for launching a new wave of fraud.
@@
To tackle these issues, some poisoning-based approaches [24; 26] have been recently proposed to perturb user images in an imperceptible way to render them “unlearnable” from malicious uses. Specifically, these approaches aim to craft perturbations that can mislead the personalized diffusion model generation process to compromise the generation ability. For instance, PhotoGuard [43] proposes to attack the VAE encoder or UNet decoder for crafting perturbation that aligns the perturbed latent code or denoised image to the ones of a dummy target. Similarly, Liang et al. [26] proposes AdvDM to craft protected images with a pre-trained fixed surrogate model with adversarial perturbation. Its latter version, Mist [25], considers an additional targeting loss for degrading the texture quality. Targeting art style transferring, Glaze [45] proposes to minimize the feature distance between the perturbed image and a targeting image with a pre-trained feature extractor. However, these works all focus on attacking text-to-image synthesis methods that leverage fixed diffusion models, where protection is much easier due to the inevitable adversarial vulnerability of the DNNs.
@@
To answer this question, we propose MetaCloak, a more effective and robust poisoning-based data protection approach against unauthorized subject-driven text-to-image diffusion-based synthesis. To tackle the first challenge, as well as to improve perturbation transferability, MetaCloak leverages meta-learning to learn model-agnostic perturbation over a pool of steps-staggered surrogates. With this meta-learning framework, we found that the perturbations are more effective than the previous approaches under the training setup without data transformations. To address the second challenge, i.e., improving the robustness of perturbation under the training setting with data transformation, we incorporate a transformation sampling process to craft transformation-robust perturbation. Besides, designing generation loss to guide poisoning is tricky since leveraging the “ground-truth” metrics might lead to an over-fitting problem. To address this problem, we design a denoising-error maximization loss to encourage the added perturbation to cause serious semantic distortion from the model perception, which aims to introduce “meaningless” and “hard-to-understand” patterns that can trick the model into learning false correlations. On the VGGFace2 and CelebA-HQ datasets, we show that our method significantly outperforms existing approaches under data transformations. Notably, MetaCloak can successfully fool an online personalized diffusion model training service, Replicate, in a black-box manner, demonstrating the effectiveness of MetaCloak in real-world scenarios. Our main contributions are summarized as follows.
@@ 2 Related Wroks
Avability Attack. Availability attack aims to degrade data availability for model training by injecting imperceptible perturbations into the training data. The underlying mechanism of this attack is rooted in the shortcut learning preference of deep neural networks (DNNs), where models turn to learn easy but brittle features for prediction. Existing works on avability attacks mainly focus on image classification [16; 17; 9; 53]. Unlearnable Example [16] and L2C [9] are two of the pioneers in this field, and later research Tao et al. [48] indicates that noise can be easily removed with adversarial training and data transformations. To address this concern, many works have been proposed, with techniques of min-max optimization [11], filter-based perturbation [41], and score guiding [8]. To better understand the defense mechanism, Tao et al. [48] propose frameworks for showing some certificated purification against such attacks. To adapt to other models and applications, studies are also conducted in natural language processing [18], graph learning [27], and contrastive learning [35]. To the best of our knowledge, we are the first to study robust availability attacks to personalized generation with diffusion models.
@@
Protection Against Unauthorized Subject-driven AI Synthesis. Unauthorized subject-driven AI synthesis including style transfer [21], personalized generation [40], and image inpainting [2], pose a serious threat to the privacy of individuals and art copyright [4; 50]. To address this concern, recent works [43; 55; 45; 24] propose to protect the data with poisoning-based approaches, which modify the data in an imperceptible way while causing severe degradation in the generation performance. PhotoGuard [43] proposes to attack the VAE encoder or UNet decoder for crafting perturbation that aligns the perturbed latent code or denoised image to the ones of the dummy target. Targeting degrading art style transferring, Glaze [45] proposes to minimize similar targeting loss with a pre-trained style-transfer model, with learned perceptual similarity as a penalty. Similarly, AdvDM [55] proposes to craft perturbation that minimizes the likelihood of the perturbed image with a pre-trained diffusion model, which equivalently maximizes the denoising loss during training. Its later version, Mist [25], further incorporates a texture targeting loss for a more robust and sharp pattern into the perturbed images. However, these works all focus on attacking the synthesis methods that leverage fixed diffusion models by exploiting the adversarial vulnerability of the DNNs. In contrast, our work focuses on conducting more challenging data protection against fine-tuning-based synthesis methods [40; 33]. To the best of our knowledge, Anti-DreamBooth [24] is the only work that studies data protection against fine-tuning-based synthesis approaches. However, it crafts sub-optimal perturbation and is also not robust against transformation purification. In our work, targetting fine-tuning-based synthesis methods, we study poisoning-based data protection with better effectiveness and robustness against transformations.
@@ 3 Preliminary
where $\mathbf{x}_{0}^{\prime}=\mathbf{x}$, $\text{sign}(\cdot)$ is the sign function, $i$ is the step index, and the step size $\alpha>0$.
During this generation process, the adversarial examples gradually progress in a direction that would increase the denoising loss while maintaining imperceptible perturbations. Recent works [24; 25] have demonstrated that perturbed images from this attack can effectively deceive text-to-image generation models [12; 40] to produce low-quality images.
@@ 5.1 Learning to Learn Transferable and Model-agnostic Perturbation
$X_{p}^{i+1}=\Pi_{B_{\infty}(X_{p}^{0},r)}(X_{p}^{i}+\alpha\mathrm{sign}(\nabla_{X_{p}^{i}}\mathcal{L}_{\text{gen}}^{*}(X_{\text{ref}};\hat{\mathbf{x}}_{\theta^{\prime}_{i,K}},X_{p}^{i})).$ (7)
@@ 5.2 Transformation-robust Semantic Distortion with Denoising-error Maximization
Our empirical observation indicates that the maximization of this loss can result in chaotic content and scattered texture in the generated images. Instead of introducing an additional targeting loss term [26; 25], we found that our simple denoising-maximization loss is more effective against fine-tuning-based diffusion models. However, perturbations crafted directly with (9) are fragile to minor transformations and ineffective in bypassing (5). To remedy this, we adopt the expectation over transformation technique (EOT [1]) into the PGD process. Specifically, given $T$ as a distribution over a set of transformations in (5), we apply EOT on (7) as
$X_{p}^{i+1}=\mathbb{E}_{g\sim T}\left[\Pi_{B_{\infty}(X_{p}^{0},r)}(X_{p}^{i}+\alpha\mathrm{sign}(\nabla_{X_{p}^{i}}\mathcal{L}_{\text{gen}}(g(X_{p}^{i});\hat{\mathbf{x}}_{\theta^{\prime}_{i,K}}))\right],$ (10)
@@ 6.1 Setup
Datasets. Our experiments are performed on human subjects using the two face datasets: Celeba-HQ [20] and VGGFace2 [3] following Anti-DreamBooth [24]. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of 30,000 celebrity face images. VGGFace2 is a comprehensive dataset with over 3.3 million face images from 9,131 unique identities. Fifty identities are selected from each dataset, and we randomly pick 8 images from each individual and split those images into two subsets for image protection and reference. More results on non-face data are in the App. C.3.
@@
Baselines and Implementation Details. We compare our method with the following adopted state-of-the-art baselines in [26; 24; 25]: i) ASPL [24] alternatively update the perturbations and surrogate models, where the surrogate models are updated on both poisons and clean data; ii) E-ASPL is an extension of ASPL that ensembles multiple types of diffusion models for better transferability; iii) FSMG leverages a DreamBooth trained on clean image for crafting adversarial examples; iv) AdvDM [26; 25] leverages a pre-trained diffusion model for crafting adversarial examples with additional targeting loss for texture distortion; v) Glaze [45] minimizes the representation between the perturbed image and a target dummy image with a pre-trianed encoder; vi) PhotoGuard [43] perturb image to align its denoised image closer to a dummy target with efficient partial back-propagation; Following the common setup [24], we set the noise radius ($\ell_{\infty}$-norm ball) to 11/255 with a step size of 1/255 and a step number of 6 by default. We set the unrolling number $K=1$, surrogate model number $M=5$, sample batch size $B=1$, and crafting step $C=4000$ with SD v2-1-base as the surrogate model, which takes about 3 GPU hours to train instance-wise noise. See App. A for more details.
@@ 6.2 Effectiveness of MetaCloak
Effectiveness Comparison through Quantitative and Qualitative Metrics. As observed in Tab. 1, MetaCloak consistently outperforms other baselines across all the metrics. Specifically, in the most important metric, SDS, which measures whether a face appeared in the generated image, MetaCloak successfully degraded this metric by 36.9% and 45.0% compared to previous SOTA on VGGFace2 and CelebA-HQ. Regarding reference-based semantic matching metrics, the results on IMS-VGG and IMS-CLIP also show that our method is more effective than other baselines. Regarding image quality metrics, the results on CLIP-IQAC suggest that MetaCloak can effectively degrade the image quality of generated images. Furthermore, the results in Tab. 5 (see appendix) also suggest that MetaCloak also achieves better effectiveness than existing methods under Stand. Training settings. For visualization, as we can see in Fig. 2, compared to other baselines, MetaCloak can robustly fool the model to generate images with low quality and semantic distortion under both Stand. and Trans. training while others are sensitive to transformations. More visualizations are in App C.2.
@@
Effectiveness under Online Training-as-services Scenarios. To test the effectiveness of our framework in the wild, we conduct experiments under online training-as-service settings. Unlike local training, attacking online training services is more challenging due to the limited knowledge of data prepossessing. We first showcase the performance of our method under two common DreamBooth fine-tuning scenarios, including full fine-tuning (Full-FT) and LoRA-fine-tuning (LoRA-FT). We sample data from VGGFace2 and upload its clean and poisoned images to Replicate [36] for DreamBooth training. From the results in Fig. 3, we can see that MetaCloak performs well under the Full-FT setting; for instance, it successfully degrades the SDS from 98.9% to 21.8%. Under the LoRA-FT setting, the model seems to be more resilient, but MetaCloak can still cause some degradation. Additionally, under another popular training setting with the Text Inversion, we validate that MetaCloak can still work well in disturbing the generation quality in Tab. 2. These results demonstrate that MetaCloak can seriously threaten online training services.
@@
Effectiveness across Different Training Settings. While crafting specifically with the SD v2-1-base model, one might wonder whether our protection can transfer across training settings with other diffusion models. To investigate that, we further conducted experiments under training settings with two other versions of stable diffusion models, including SD v2-1 and SD v1-5. The results in Fig. 5 show that our protection is well transferable in introducing distortion in generation across different versions of stable diffusion models. Please refer to the App. C.4 for more results. Furthermore, we also study training settings with different protection ratios and radii under the Trans. Training settings on VGGFace2 dataset. As shown in Fig. 4, these two factors are essential for protection. Besides, MetaCloak is more budget-efficient since it matches the same performance as the previous SOTA with lower protection ratios and radii. Please refer to the App. C.7 on the trade-off between stealthiness and performance of our perturbation under different radii.
@@ 6.3 Resistance against Adversarial Purification
We consider three adversarial purification techniques, including JPEG compression [28], super-resolution transformation (SR) [31], and image reconstruction based on total-variation minimization (TVM) [51]. We use a quality factor of 75 for the JPEG defense and a scale factor of 4 for the SR defense [26]. As shown in Tab. 4, additional purification does hinder the data protection performance of MetaCloak to some extent, and SR is the most effective one in restoring the generation ability compared to others. For the other two approaches, we found that JPEG compression will compromise the purified image quality, and the TVM does better in retaining graphical quality, but it causes significant and unacceptable distortion in the face region of purified images. While SR does the best in purification, it still can not fully recover the original generation ability of DreamBooth trained on clean data. See App. C.6 for more details and results.

### v3

# MetaCloak: Preventing Unauthorized Subject-driven Text-to-image Diffusion-based Synthesis via Meta-learning

@@ 1 Introduction
Diffusion models achieve significant success in a wide range of applications, including image generation [16, 51, 8], image editing [23, 50, 6], and text-to-image synthesis [42]. Subject-driven text-to-image synthesis, an emerging application of diffusion models, in particular, has attracted considerable attention due to its potential to generate personalized images from a few reference photos. Among the approaches proposed to achieve this goal [38, 46], DreamBooth [44] and Text Inversion [13] are two prominent training-based methods that offer impressive personalize generation ability. With an additional lightweight model fine-tuning process for capturing the subject or a concept-related embedding training phase, personalized diffusion models can retain the generation capacity from the pre-training stage and conduct vivid personalized generation. While these methods empower high-quality personalized generation, they also raise privacy concerns as they can fabricate misleading or harmful content in the wrong hands, endangering individuals. For example, recent news [35, 3, 21] indicates that AI tools like diffusion models have been employed to generate fake profiles of individuals for launching a new wave of fraud.
@@
To tackle these issues, some poisoning-based approaches [26, 28] have been recently proposed to perturb user images in an imperceptible way to render them “unlearnable” from malicious uses. Specifically, these approaches aim to craft perturbations that can mislead the personalized diffusion model generation process to compromise the generation ability. For instance, PhotoGuard [47] proposes to attack the VAE encoder or UNet decoder for crafting perturbation that aligns the perturbed latent code or denoised image to the ones of a dummy target. Similarly, Liang et al. [28] proposes AdvDM to craft protected images with a pre-trained fixed surrogate model with adversarial perturbation. Its latter version, Mist [27], considers an additional targeting loss for degrading the texture quality. Targeting art style transferring, Glaze [49] proposes to minimize the feature distance between the perturbed image and a targeting image with a pre-trained feature extractor. However, these works all focus on attacking text-to-image synthesis methods that leverage fixed diffusion models, where protection is much easier due to the inevitable adversarial vulnerability of the DNNs.
@@
To answer this question, we propose MetaCloak, a more effective and robust poisoning-based data protection approach against unauthorized subject-driven text-to-image diffusion-based synthesis. To tackle the first challenge, as well as to improve perturbation transferability, MetaCloak leverages meta-learning to learn model-agnostic perturbation over a pool of steps-staggered surrogates. With this meta-learning framework, we found that the perturbations are more effective than the previous approaches under the training setup without data transformations. To address the second challenge, i.e., improving the robustness of perturbation under the training setting with data transformation, we incorporate a transformation sampling process to craft transformation-robust perturbation. Besides, designing generation loss to guide poisoning is tricky since leveraging the “ground-truth” metrics might lead to an over-fitting problem. To address this problem, we design a denoising-error maximization loss to encourage the added perturbation to cause serious semantic distortion from the model perception, which aims to introduce “meaningless” and “hard-to-understand” patterns that can trick the model into learning false correlations. On the VGGFace2 and CelebA-HQ datasets, we show that our method significantly outperforms existing approaches under data transformations. Notably, MetaCloak can successfully fool online personalized diffusion model training services like Replicate in a black-box manner, demonstrating the effectiveness of MetaCloak in real-world scenarios. Our main contributions are summarized as follows.
@@ 2 Related Wroks
Availability Attack. Availability attack aims to degrade data availability for model training by injecting imperceptible perturbations into the training data. The underlying mechanism of this attack is rooted in the shortcut learning preference of deep neural networks (DNNs), where models turn to learn easy but brittle features for prediction. Existing works on avability attacks mainly focus on image classification [17, 18, 10, 57]. Unlearnable Example [17] and L2C [10] are two of the pioneers in this field, and later research Tao et al. [52] indicates that noise can be easily removed with adversarial training and data transformations. To address this concern, many works have been proposed, with techniques of min-max optimization [12], filter-based perturbation [45], and score guiding [9]. To better understand the defense mechanism, Tao et al. [52] propose frameworks for showing some certificated purification against such attacks. To adapt to other models and applications, studies are also conducted in natural language processing [20], graph learning [29], and contrastive learning [39]. To the best of our knowledge, we are the first to study robust availability attacks to personalized generation with diffusion models.
@@
Protection Against Unauthorized Subject-driven AI Synthesis. Unauthorized subject-driven AI synthesis including style transfer [23], personalized generation [44], and image inpainting [2], pose a serious threat to the privacy of individuals and art copyright [5, 54]. To address this concern, recent works [47, 59, 49, 26] propose to protect the data with poisoning-based approaches, which modify the data in an imperceptible way while causing severe degradation in the generation performance. PhotoGuard [47] proposes to attack the VAE encoder or UNet decoder for crafting perturbation that aligns the perturbed latent code or denoised image to the ones of the dummy target. Targeting degrading art style transferring, Glaze [49] proposes to minimize similar targeting loss with a pre-trained style-transfer model, with learned perceptual similarity as a penalty. Similarly, AdvDM [59] proposes to craft perturbation that minimizes the likelihood of the perturbed image with a pre-trained diffusion model, which equivalently maximizes the denoising loss during training. Its later version, Mist [27], further incorporates a texture targeting loss for a more robust and sharp pattern into the perturbed images. However, these works all focus on attacking the synthesis methods that leverage fixed diffusion models by exploiting the adversarial vulnerability of the DNNs. In contrast, our work focuses on conducting more challenging data protection against fine-tuning-based synthesis methods [44, 37]. To the best of our knowledge, Anti-DreamBooth [26] is the only work that studies data protection against fine-tuning-based synthesis approaches. However, its protection is brittle under transformation purification. In this work, we study poisoning-based data protection with better effectiveness and robustness against transformations.
@@ 3 Preliminary
where $\mathbf{x}_{0}^{\prime}=\mathbf{x}$, $\text{sign}(\cdot)$ is the sign function, $i$ is the step index, and the step size $\alpha>0$. During this generation process, the adversarial examples gradually progress in a direction that would increase the denoising loss while maintaining imperceptible perturbations. Recent works [26, 27] have demonstrated that perturbed images from this attack can effectively deceive text-to-image generation models [13, 44] to produce low-quality images.
@@ 5.1 Learning to Learn Transferable and Model-agnostic Perturbation
$\scriptsize X_{p}^{i+1}=\Pi_{B_{\infty}(X_{p}^{0},r)}(X_{p}^{i}+\alpha\mathrm{sign}(\nabla_{X_{p}^{i}}\mathcal{L}_{\text{gen}}^{*}(X_{\text{ref}};\hat{\mathbf{x}}_{\theta^{\prime}_{i,K}},X_{p}^{i})).$ (7)
@@ 5.2 Transformation-robust Semantic Distortion with Denoising-error Maximization
Our empirical observation indicates that the maximization of this loss can result in chaotic content and scattered texture in the generated images. Instead of introducing an additional targeting loss term [28, 27], we found that our simple denoising-maximization loss is more effective against fine-tuning-based diffusion models. However, perturbations crafted directly with (9) are fragile to minor transformations and ineffective in bypassing (5). To remedy this, we adopt the expectation over transformation technique (EOT [1]) into the PGD process. Specifically, given $T$ as a distribution over a set of transformations in (5), we apply EOT on (7) as
$\scriptsize X_{p}^{i+1}=\mathbb{E}_{g\sim T}\left[\Pi_{B_{\infty}(X_{p}^{0},r)}(X_{p}^{i}+\alpha\mathrm{sign}(\nabla_{X_{p}^{i}}\mathcal{L}_{\text{gen}}(g(X_{p}^{i});\hat{\mathbf{x}}_{\theta^{\prime}_{i,K}}))\right],$ (10)
@@ 6.1 Setup
Datasets. Our experiments are performed on human subjects using the two face datasets: Celeba-HQ [22] and VGGFace2 [4] following Anti-DreamBooth [26]. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of 30,000 celebrity face images. VGGFace2 is a comprehensive dataset with over 3.3 million face images from 9,131 unique identities. Fifty identities are selected from each dataset, and we randomly pick eight images from each individual and split those images into two subsets for image protection and reference. More results on non-face data are in the App. C.3.
@@
Baselines and Implementation Details. We compare our method with the following adopted state-of-the-art baselines in [28, 26, 27]: i) ASPL [26] alternatively update the perturbations and surrogate models, where the surrogate models are updated on both poisons and clean data; ii) E-ASPL is an extension of ASPL that ensembles multiple types of diffusion models for better transferability; iii) FSMG leverages a DreamBooth trained on clean image for crafting adversarial examples; iv) AdvDM [28, 27] leverages a pre-trained diffusion model for crafting adversarial examples with additional targeting loss for texture distortion; v) Glaze [49] minimizes the representation between the perturbed image and a target dummy image with a pre-trianed encoder; vi) PhotoGuard [47] perturb image to align its denoised image closer to a dummy target with efficient partial back-propagation; Following the common setup [26], we set the noise radius ($\ell_{\infty}$-norm ball) to 11/255 with a step size of 1/255 and a step number of 6 by default. We set the unrolling number $K=1$, surrogate model number $M=5$, sample batch size $B=1$, and crafting step $C=4000$ with SD v2-1-base as the surrogate model, which takes about 3 GPU hours to train instance-wise noise. See App. A for more details.
@@ 6.2 Effectiveness of MetaCloak
Effectiveness Comparison through Quantitative and Qualitative Metrics. As observed in Tab. 1, MetaCloak consistently outperforms other baselines across all the metrics. Specifically, in the most important metric, SDS, which measures whether a face appeared in the generated image, MetaCloak successfully degraded this metric by 36.9% and 45.0% compared to previous SoTA on VGGFace2 and CelebA-HQ. Regarding reference-based semantic matching metrics, the results on IMS-VGG and IMS-CLIP also show that our method is more effective than other baselines. Regarding image quality metrics, the results on CLIP-IQAC suggest that MetaCloak can effectively degrade the image quality of generated images. Furthermore, the results in Tab. 5 (see App. C.1) also suggest that MetaCloak also achieves better effectiveness than existing methods under Stand. Training settings. For visualization, as we can see in Fig. 2, compared to other baselines, MetaCloak can robustly fool the model to generate images with low quality and semantic distortion under both Stand. and Trans. training while others are sensitive to transformations. More visualizations are in App C.2.
@@
Effectiveness under Online Training-as-services Scenarios. To test the effectiveness of our framework in the wild, we conduct experiments under online training-as-service settings. Unlike local training, attacking online training services is more challenging due to the limited knowledge of data prepossessing. We first showcase the performance of our method under two common DreamBooth fine-tuning scenarios, including full fine-tuning (Full-FT) and LoRA-fine-tuning (LoRA-FT). We sample data from VGGFace2 and upload its clean and poisoned images to Replicate [40] for DreamBooth training. From the results in Fig. 3, we can see that MetaCloak performs well under the Full-FT setting; for instance, it successfully degrades the SDS from 98.9% to 21.8%. Under the LoRA-FT setting, the model seems to be more resilient, but MetaCloak can still cause some degradation. On the inherent defense ability of LoRA-FT, we conjecture that it only fine-tunes a few additional layers of parameters, which might be less likely to overfit and thus more robust to MetaCloak. Additionally, under another popular training setting with the Text Inversion, we validate that MetaCloak can still work well in disturbing the generation quality in Tab. 2. These results demonstrate that MetaCloak can seriously threaten online training services.
@@
Effectiveness across Different Training Settings. While crafting specifically with the SD v2-1-base model, one might wonder whether our protection can transfer across training settings with other diffusion models. To investigate that, we further conducted experiments under training settings with two other versions of stable diffusion models, including SD v2-1 and SD v1-5. The results in Fig. 4 show that our protection is well transferable in introducing distortion in generation across different versions of stable diffusion models. Furthermore, we also study training settings with different protection ratios and radii under the Trans. Training settings on VGGFace2 dataset. As shown in Fig. 5, these two factors are essential for protection. Besides, MetaCloak is more budget-efficient since it matches the performance of the previous SoTA with lower protection ratios and radii. Please refer to the App. C.6 on the trade-off between stealthiness and performance of our perturbation under different radii.
@@ 6.3 Resistance against Adversarial Purification
We consider three adversarial purification techniques, including JPEG compression [30], super-resolution (SR) [33], and image reconstruction based on total-variation minimization (TVM) [55]. We use a quality factor of 75 for the JPEG defense and a scale factor of 4 for the SR defense [28]. As shown in Tab. 7, additional purification does hinder the data protection performance of MetaCloak to some extent, and SR is the most effective one in restoring the generation ability compared to others. For the other two approaches, we found that JPEG compression will compromise the purified image quality, and the TVM does better in retaining graphical quality, but it causes significant and unacceptable distortion in the face region of purified images. While SR does the best in purification, it still can not fully recover the original generation ability of DreamBooth trained on clean data. See App. C.5 for more details and results.